### v2

# Effects of Lorentz invariance violation on particle and photon production in astrophysical sources

Abstract
We investigate the impact of Lorentz invariance violation (LIV) on radiation processes in astrophysical sources, focusing on synchrotron and inverse Compton interactions. We derive modified expressions for radiated power and photon energy under LIV assumptions and incorporate them into first-order Fermi acceleration models. Our analysis reveals energy thresholds beyond which LIV significantly alters particle dynamics and photon spectra, introducing non-physical divergences that highlight limitations in perturbative approaches. We model synchrotron self-Compton (SSC) emission in the presence of LIV and assess its consequences for photon fluxes from blazars, including Markarian 501 and the BL Lac population. LIV introduces distinct high-energy emission regions that deviate from standard expectations. Comparisons with observational data, particularly upper limits from the Pierre Auger Observatory, suggest that future multi-messenger observations could constrain LIV parameters through the non-detection of such excesses.
1 Introduction
Lorentz invariance (LI) is a fundamental principle in modern physics, underlying the theory of relativity [1] and the standard model of particle physics [2, 3]. It asserts that the laws of physics remain the same across all inertial frames of reference. However, potential violations of Lorentz invariance (LIV) [4] have been theorized within the context of quantum gravity [5, 6, 7] and other theories beyond the standard model [8, 9]. Investigating these possible violations provides insight into fundamental aspects of the Universe, potentially revealing new physics at extremely high energies.
The propagation of astroparticles in the Universe [10, 11, 12, 13, 14] and in the development of extensive air showers [15, 16] have imposed limits on LIV parameters [17, 18, 19, 20, 21]. Nonetheless, a thorough investigation of the processes leading to the production of charged particles and photons under the hypothesis of LIV remains necessary [22].
The mechanisms generating these particles and photons are fundamentally dependent on LI, which plays a critical role in determining the maximum energy, the emitted power, and the resulting energy spectrum [23]. Consequently, such sources provide a natural framework for probing potential modifications of LI. In particular, the detection of highly energetic astroparticles accompanied by photon counterparts offers a promising framework for probing potential deviations from LI within the context of multi-messenger astrophysics.
In this work, we investigate how LIV changes synchrotron and inverse Compton [24] interactions, the two most important energy loss processes in astrophysical sources [25, 26, 27, 28, 29, 30, 31]. We also incorporate the LIV models for synchrotron and inverse Compton interactions into the context of first-order Fermi acceleration and photon emission in blazars. We investigate how the production of high-energetic particles and photons is affected by the LIV assumption.
In Section 2, we present the LIV framework. In Subsection 3.1, we derive the expressions for the LIV-modified synchrotron radiated power and the corresponding synchrotron photon energy. The application of LIV to the inverse Compton scattering process is addressed in Subsection 3.2. A combined treatment of synchrotron and inverse Compton emission within a synchrotron self-Compton (SSC) scenario, incorporating LIV effects, is developed in Section 4, where we analyze a case study of first-order Fermi acceleration. In this context, we independently evaluated the impact of LIV on both energy losses and energy spectrum of particles, as well as on the resulting photon emission. Section 5 presents a comparison between our theoretical predictions and observational data, including the identified blazar population and upper limits on photon flux from the Pierre Auger Observatory. Finally, our conclusions and implications are summarized in Section 6.
2 LIV framework
Lorentz Invariance Violation (LIV), as predicted by various theoretical models [32, 33, 34], leads to a modification of Einstein’s energy-momentum dispersion relation. This deviation can be expressed by additional terms in the standard dispersion relation:
$E^{2}=m^{2}+p^{2}+\sum_{n}\delta_{n}p^{n+2}\rm{,}$ (2.1)
where we have set $c=1$. The momentum coefficients $\delta_{n}$ represent the parameters responsible for the strength of LIV, while $n$ denotes the order of violation. Considering that we focus on highly energetic particles, we assume $p\gg m$, which leads to
$p=\frac{E}{\sqrt{1+\delta_{n}E^{n}}}\rm{.}$ (2.2)
Lorentz Invariance Violation will also modify the Lorentz factor and, as shown in [35], this shall be done via the group velocity. Considering the definition $\gamma^{2}=\left(1-v^{2}\right)^{-1}$, as well as the group velocity, $v=\frac{\partial E}{\partial p}$, the LIV-modified Lorentz factor results in
$\gamma^{2}_{\textrm{LIV}}=\frac{E^{2}}{{m^{2}-\left(n+1\right)\delta_{n}E^{n+2}}}\textrm{,}$ (2.3)
where the high-energy approximation was used. As can be observed, the Lorentz factor increases even faster with energy under the LIV assumption and tends to infinity when $E^{n+2}\rightarrow\frac{m^{2}}{(n+1)\delta_{n}}$. Once this occurs, the Lorentz factor becomes negative, revealing a fundamental limitation of the current model and indicating the need for a non-perturbative treatment.
Different particles may exhibit different breaking parameters [4]. Therefore, we will denote the breaking parameters for electrons as $\delta_{n}^{(e)}$, for protons as $\delta_{n}^{(p)}$ and for photons as $\delta_{n}^{(\gamma)}$. When $\delta_{n}$ is used without specifying the type of particle, a charged particle is being considered.
3 Radiation processes in particle acceleration under LIV assumption
Accelerated charged particles emit electromagnetic radiation as described by Larmor’s formula [23, 36, 37]. The emitted power is
$P=\frac{2}{3}q^{2}\gamma^{6}\left[|\dot{\vec{v}}|^{2}-|\vec{v}\times\dot{\vec{v}}|\right]\rm{,}$ (3.1)
where $q$ is the charge of the particle, $v$ and $\gamma$ are, respectively, its velocity and its Lorentz factor in the laboratory frame of reference. In the next subsections, we solve synchrotron and inverse Compton radiation considering LIV.
3.1 Synchrotron radiation under LIV assumption
A charged particle moving with velocity $\vec{v}$ through a magnetic field ($\vec{B}$) is subject to the Lorentz force, $\vec{F}=q\vec{v}\times\vec{B}$. This acceleration results in the emission of electromagnetic radiation, a process known as synchrotron radiation. Given that the force is always perpendicular to the velocity $\vec{F}=q\vec{v}\times\vec{B}=\gamma m\dot{\vec{v}}$, the power emitted in synchrotron radiation is
$P=\frac{2}{3}\frac{q^{4}}{m^{2}}\gamma^{2}|\vec{v}\times\vec{B}|^{2}\,.$ (3.2)
The energy of the emitted photons is highly concentrated at the peak of the spectral power distribution [38]. For simplicity, we assume the charged particle emits all photons with the same energy at the peak frequency
$E_{\gamma}=0.29\cdot\frac{3}{2}\gamma^{2}\frac{qB}{m}\,.$ (3.3)
Using equation 3.4
$P^{\text{LIV}}=\frac{2}{3}\frac{q^{4}}{m^{4}}|\vec{v}\times\vec{B}|^{2}\left[\frac{E^{2}}{1-\left(n+1\right)\frac{\delta_{n}}{m^{2}}E^{n+2}}\right]$ (3.4)
and
$E_{\gamma}^{\text{LIV}}=0.29\cdot\frac{3}{2}\cdot\frac{qB}{m^{3}}\left[\frac{E^{2}}{1-\left(n+1\right)\frac{\delta_{n}}{m^{2}}E^{n+2}}\right]\rm{.}$ (3.5)
LIV has a significant effect on the synchrotron radiation. The modification of the Lorentz factor leads to a discontinuity at $E^{n+2}=\frac{m^{2}}{(n+1)\delta_{n}}$ in the power emitted and energy of the emitted photons. For particles with energy $E^{n+2}>\frac{m^{2}}{(n+1)\delta_{n}}$, the energy of the emitted photon becomes negative. These consequence of the calculations indicate a limitation in the perturbative approach used here. Introducing new physics for scales beyond the breaking point may be necessary to fully describe such events. Similar consequences were obtained by different approaches [39, 40].
3.2 Inverse Compton under LIV assumption
Inverse Compton scattering is the interaction in which a low-energy photon gains energy by scattering off a high-energy charged particle. The energy of the scattered photon can be calculated imposing the conservation of the Lorentz-invariant 4-momentum
$p_{p}^{\mu}+p_{\gamma}^{\mu}=p_{p}^{{}^{\prime}\mu}+p_{\gamma}^{{}^{\prime}\mu}\textrm{,}$ (3.6)
where $p$ denotes the charged particle, $\gamma$ denotes photon and primed quantities refer to final states.
The scattering cross section of relativistic particles is given by the Klein-Nishina formula [24]
[…]
and the average scattering angle can be calculated through
[…]
Introducing LIV via the energy-momentum in Equation 2.1, the spatial components ($\mu=1,2,3$) become
$E^{{}^{\prime}2}_{p}=m^{2}+\left(\frac{E_{\gamma}^{{}^{\prime}2}}{1+\delta_{n}^{(\gamma)}E_{\gamma}^{{}^{\prime}n}}+\frac{E_{\gamma}^{2}}{1+\delta_{n}^{(\gamma)}E_{\gamma}^{n}}-\frac{2E_{\gamma}E_{\gamma}^{{}^{\prime}}\cos{\theta}}{\sqrt{\left(1+\delta_{n}^{(\gamma)}E_{\gamma}^{{}^{\prime}n}\right)\left(1+\delta_{n}^{(\gamma)}E_{\gamma}^{n}\right)}}\right)\times\\ \times\left[1+\delta_{n}\left(\frac{E_{\gamma}^{{}^{\prime}2}}{1+\delta_{n}^{(\gamma)}E_{\gamma}^{{}^{\prime}n}}+\frac{E_{\gamma}^{2}}{1+\delta_{n}^{(\gamma)}E_{\gamma}^{n}}-\frac{2E_{\gamma}E_{\gamma}^{{}^{\prime}}\cos{\theta}}{\sqrt{\left(1+\delta_{n}^{(\gamma)}E_{\gamma}^{{}^{\prime}n}\right)\left(1+\delta_{n}^{(\gamma)}E_{\gamma}^{n}\right)}}\right)^{n/2}\right]^{1/2}\textrm{,}$ (3.9)
where $\theta$ is the scattering angle, while the time component ($\mu=0$) yields
$m^{2}+2mE_{\gamma}+E^{2}_{\gamma}-E^{{}^{\prime}2}_{\gamma}-E_{p}^{{}^{\prime}2}-2E^{\prime}_{\gamma}E^{\prime}_{p}=0\,\textrm{.}$ (3.10)
Reference [41] showed similar results for the case $n=1$. However, here, we present a generalization for all values of $n$. For this interaction, LIV effects could be considered in the charged particle sector and the photon sector. In the following, we focus on the photon sector and defer the full treatment of the charged-particle sector to future work.
We set $\delta_{n}=0$ to isolate effects on the photon sector and solved Equation 3.10 numerically. Figure 1 shows the LIV effect on the energy of the photon after the interaction. The plot shows how the energy of the photon after the interaction (final photon energy) depends on the energy of the photon before the interaction (initial photon energy) for different LIV coefficients. The energy of the photon after the interaction tends to the Thomson regime when the initial photon energy is larger than a value depending on the LIV coefficient, as shown in the plot. This can be verified as Equation 3.9 and Equation 3.10 leads to $E_{\gamma}^{\prime}=E_{\gamma}$, which is the Thomson regime, in the high-energy limit ($\delta_{n}^{(\gamma)}E_{\gamma}^{n}\gg 1$).
Figure 2 shows the evolution of the cross section with the initial energy of the photon, including LIV via Equations 3.9 and 3.10. LIV significantly impacts $\sigma_{\textrm{KN}}$, enhancing the interaction probability and effectively bypassing Klein-Nishina suppression until the Thomson limit is recovered.
Figure 3 shows the evolution of the scattering angle with the initial energy of the photon including LIV. A new transition to the Thomson regime is observed at high energies.
Using the modified cross section and final energy of the photon, the emitted power under LIV assumption can be calculated via
[…]
where $\eta(E_{\gamma})$ is the energy distribution of photons before the interaction. LIV increases both the final energy and cross section, thus amplifying the emitted power.
4 First-order Fermi acceleration including radiation process under LIV assumption
First-order Fermi mechanism [42, 43, 44] is an efficient acceleration process known to be working in several astrophysical sources. In a previous publication [22], we investigated how LIV modifies the energy gain per scattering, the acceleration timescales, and the spectral indices of particles accelerated by the first-order Fermi mechanism. In this Section, we will merge the first-order Fermi with synchrotron and SSC energy losses, both including LIV.
4.1 Synchrotron radiation in first-order Fermi acceleration under LIV assumption
The synchrotron loss process can be highly relevant in scenarios of astrophysical particle acceleration, mainly in two stages: initially, during the acceleration of the charged particles itself, where energy gains compete with radiative losses, and, subsequently, near the source, where the particles interact with the radiation fields.
4.1.1 Synchrotron: maximum energy of accelerated particles
In Reference [22], we calculated the energy gained by particles under first-order Fermi acceleration considering LIV. This modified energy gain, along with the power lost through synchrotron radiation, gives the following net gain $-b$:
$b=-\frac{\left<\Delta E\right>}{\tau_{avg}}+P_{sync}\frac{\tau_{avg}}{\tau_{avg}}\rm{,}$ (4.1)
where $\tau_{avg}$ is the average crossing time for the particle traversing the shock wave. We consider the best-case scenario of Bohm diffusion [45]
$\tau_{avg}\approx\frac{5}{qBV}\frac{E}{\sqrt{1+\delta_{n}E^{n}}}\rm{.}$ (4.2)
In a scenario where synchrotron radiation is the dominant energy loss during shock wave acceleration, we can use Equation 3.4 to determine the maximum energy a charged particle can gain through first-order Fermi acceleration [46] as the equilibrium between loss and energy gain, $\Delta E\right>=P_{sync}\cdot\tau_{avg}$, which gives
$\frac{E_{\textrm{max}}^{2}}{\left[1-\left(n+1\right)\frac{\delta_{n}}{m^{2}}E^{n+2}\right]}\approx\frac{m^{4}}{B}\rm{.}$ (4.3)
Figure 4 and Figure 5 show the solution to this equation as a function of $\delta_{n}$ for electrons and protons, respectively. A transition between the two regimes is very clear. In the regime of low $\delta_{n}$ values, the maximum energy is determined by the balance between energy gain and synchrotron losses. As $\delta_{n}$ increases, LIV becomes more significant, the Lorentz factor barrier dominates, such that a particle reaching $E^{n+2}=\frac{m^{2}}{(n+1)\delta_{n}}$ would emit all its energy, as shown in Equation 3.4. This transition is shown as an abrupt energy loss in these plots.
4.1.2 Synchrotron: energy spectrum of particles and photons
We will use the one-zone model to calculate the energy spectrum of particles accelerated in the source [47]. This model considers particles traveling through an acceleration and a cooling (or energy-loss) region. The transport of particles can be calculated using the diffusion-loss equation to describe the number of particles $N$ within an energy range $E$ to $E+\delta E$ at a given time
$\frac{dN}{dt}+\frac{d}{dE}\left[b\cdot N\right]=-\frac{N}{\tau_{scp}}+Q_{\mathrm{inj}}\rm{,}$ (4.4)
where $b$ is the energy loss rate, in this case dominated by synchrotron radiation, $\tau_{\mathrm{esc}}$ is the escape timescale from the cooling region (here neglected), and $Q_{\mathrm{inj}}(E)$ is the injection spectrum from the acceleration site. We model $Q_{\mathrm{inj}}$ as a first-order Fermi spectrum with LIV and an exponential cutoff at $E_{\mathrm{max}}$ (derived in Subsection 4.1.1).
Figures 6 and 7 show the energy spectrum of charged particles resulting from the one-zone model, including LIV for electrons and protons, respectively. A magnetic field of $B=10^{-10}\ \textrm{T}$ was used. The suppression in the energy spectrum is a direct consequence of the barrier created by the Lorentz factor. As shown in section 4.1.1, above a certain maximum energy, which depends on the LIV break, energy losses are dominant, and particles stop gaining energy.
The energy spectrum of photons produced by the accelerated particles via synchrotron emission can be modeled by [23]
[…]
where $P$ is the synchrotron power, $N_{p}$ is the particle’s energy spectrum, and $D$ is the volume of the source. As previously discussed, we assume all photons are emitted with the same energy at the peak frequency, resulting in
$N_{\gamma}(E_{\gamma})\approx 10^{14}\frac{DB}{m}N_{p}(E)E\left[\frac{1-\left(n+1\right)\frac{\delta_{n}}{m^{2}}E^{n+2}}{1+(n+1)^{2}\frac{\delta_{n}}{m^{2}}E^{n+2}}\right]\rm{,}$ (4.6)
where $E$ is the energy of the particle that emits a photon with energy $E_{\gamma}$ according to Equation 3.5, and $N_{p}(E)$ is the spectrum calculated using Equation 4.4.
Figures 8 and 9 show the resulting energy spectrum for electrons and protons, respectively. In these calculations, we considered a source with a size of $1\hskip 3.00003pt\rm{pc}$.
LIV leads to the emergence of a distinct high-energy emission, where the final photon energy exhibits rapid growth near the divergence threshold - $E^{n+2}\rightarrow\frac{m^{2}}{(n+1)\delta_{n}}$ expected from Equation 3.5. Notably, this region contributes a non-negligible flux, consistent with the increasing emitted power derived in Equation 3.4. However, it is important to note that the energy of the emitted photon exceeds that of the electron or proton responsible for its production, thus violating energy conservation. The energy range where this violation occurs is indicated by a dashed line in Figures 8 and 9.
Lorentz invariance violation not only introduces an unanticipated energy domain in synchrotron emission but could also alter SSC spectra through the inclusion of higher-energy photons in the radiation background, resulting in emissions at previously unexpected energies and potentially altering the predicted spectra. These modifications could serve as a means to constrain LIV, as the non-detection of such new energy regions could impose limits on the LIV parameters.
4.2 Synchrotron self-Compton in first-order Fermi acceleration under LIV assumption
In non-thermal astrophysical spectra, the population of higher-energy photons is commonly attributed to the inverse Compton scattering. In this process, background photons near the source gain energy from charged particles in the acceleration region.
In regions with a high concentration of electrons and strong magnetic fields, the SSC model can become dominant [48, 49, 23]. In this framework, the seed photons originate from the synchrotron radiation produced by the particle population itself. This study will focus on this type of process, where the photon density in Equation 3.11 is determined by the synchrotron emission presented in Equation 4.6, while the cross section and final photon energy are the same as discussed in Subsection 3.2. Notice that $\delta_{n}^{(\gamma)}$ will not affect the photon distribution; only the inverse Compton scattering will be affected by this parameter. Again, the parameter $\delta_{n}$ will be set to 00 in order to focus exclusively on the behavior of LIV in photons.
4.2.1 Synchrotron self-Compton: energy spectrum of particles and photons
In this section, we repeat the calculations in Subsection 4.1.2, now considering SSC as the dominant energy-loss mechanism. Figures 10 and 11 show the energy spectrum of electrons and protons, respectively. A suppression of the flux with increasing energy is seen in both cases. The suppression caused by LIV is well below in energy and has less abrupt fall when compared to the cutoff in SSC under LI assumption. The LIV suppression is caused by two effects, the final energy in the interaction and the cross section, which tend to increase rapidly when LIV is considered, inducing a growth in the emitted power.
Following the same arguments of the previous section, we also calculated the energy spectrum of photons produced in the SSC interaction of the charged particles. Figure 12 and 13 show the energy spectrum of photons resulting from the interaction of electrons and protons, respectively. A magnetic field of $B=10^{-10}\ \textrm{T}$ was considered.
The photons emitted by protons and electrons exhibit similar behavior, differing primarily in their LIV parameters due to the proton’s higher energy. As shown in Figures 12 and 13, a new emission region emerges at higher energies due to LIV. The peak height depends on the magnetic field strength, with stronger fields producing more energetic background photons, enhancing LIV effects. The charged-particle cutoff energy also influences emission in this region — lower cutoffs reduce its prominence. We assume SSC losses dominate, significantly suppressing the particle energy spectrum (Figures 10, 11). Therefore, competition arises between Fermi-accelerated spectral shaping and emission effects: a stronger violation increases emission but suppresses the electron/proton population responsible for high-energy photons. This suppression also shifts the standard SSC peak to lower energies. In contrast, for systems dominated by synchrotron losses, only emission modifications persist without spectral suppression, as shown in Figure 14.
5 Radiation emission from blazars
In the previous sections, idealized cases of acceleration, emission, and related processes were considered to highlight the new LIV phenomena introduced in each case. In this section, we explore a more realistic scenario of photon production including LIV in Markarian 501 and in blazars.
5.1 GeV-TeV photon flux from Markarian 501
We used the models developed in the previous sections to calculate the GeV-TeV photon flux generated in Mkr 501. We used the data reference [50], in particular, the magnetic field strength and the electron energy spectrum responsible for both synchrotron emission and the inverse Compton process. We assume a cutoff at $10^{14}\ \rm{eV}$, consistent with the maximum energy predicted by Equation 4.3 under the magnetic field conditions of Mrk 501.
Figure 15 shows the energy spectrum of photons for two values of LIV break when synchrotron emission is considered according to section 3.1. The emission extends to higher energies, effectively bypassing the cutoff imposed by the LI predictions. Additionally, Figure 16, which presents a magnified view of the synchrotron peak, shows that for sufficiently large LIV parameters, even the low-energy portion of the spectrum may be altered, leading to a noticeable divergence from the observational data.
Figure 17 shows the energy spectrum of photons for two values of LIV break when SSC emission is considered according to section 3.2. A second emission region is predicted under LIV assumption which is not present in LI models. This new component in the flux allows one to open a new window to constrain the parameters associated with LIV. The non-detection of the predicted high-energy photons would impose upper limits on the extent of the violation.
5.2 Ultra-high energy photons from BL Lac
The new emission regions arising from synchrotron and SSC processes also populate the photon spectrum at energies above $10^{17}\ \rm{eV}$.
To illustrate this possibility, we compare the predicted LIV-induced emissions with the upper limits on the photon flux established by the Pierre Auger Observatory [51, 52]. The flux is modeled as originating from a population of BL Lac objects, given that approximately $10^{3}$ such sources have been identified [53, 54]. For simplicity, the combined flux is computed by summing the contributions of $10^{3}$ identical sources without accounting for source distribution, propagation effects, or intrinsic source-to-source variations, and adopting typical parameters for BL Lac objects [55]. We note that the determination of average source parameters typically does not aim to constrain the cutoff energy of the electron energy spectrum, as its impact on the overall source characteristics is minimal. Therefore, in this work we adopt a cutoff energy of $10^{14}\ \rm{eV}$.
Figures 18 and 19 show the energy spectrum of synchrotron photons calculated with this simplified model considering LIV for $n=1$ and $n=2$, respectively. We show the upper limits on the photon flux measured by the Pierre Auger Observatory for comparison in Figures 20 and 21.
A similar procedure was applied to the SSC emission considering LIV. The energy spectrum of photons are presented in Figures 22 and 23 for $n=1$ and $n=2$, respectively. Additionally, we consider a scenario without an intrinsic cutoff in the electron energy distribution, illustrating the most optimistic case for high-energy photon emission. These comparisons between the flux and the Pierre Auger Observatory limits can be seen in Figures 24 and 25.
At present, the comparison with data from the Pierre Auger Observatory does not yield constraints on the investigated LIV parameters. Establishing meaningful limits would require a more detailed astrophysical treatment, including the effects of source distribution, cosmological propagation, and potential contributions from other classes of sources, such as flat-spectrum radio quasars. Increasing the number of detected sources would enhance the cumulative photon flux, and a detailed analysis incorporating these effects could potentially lead to significant constraints on LIV.
6 Conclusion
In this work, we examined the influence of Lorentz invariance violation (LIV) on the production and emission of high-energy particles and photons in astrophysical environments. By modifying synchrotron and inverse Compton processes within the LIV framework, we derived the resulting changes in radiated power, photon energy, and energy spectra of accelerated particles. Our results demonstrate that LIV induces rapid growth in the Lorentz factor at high energies, leading to divergences in emission characteristics and suggesting a breakdown of the perturbative formalism, requiring new theoretical treatments at extreme energies.
Incorporating these effects into first-order Fermi acceleration models and synchrotron self-Compton scenarios, we showed that LIV modifies both particle spectra and the resulting multi-wavelength photon emission. In particular, LIV predicts the emergence of distinct high-energy components in the photon spectrum that are absent under Lorentz-invariant assumptions. These features were evaluated in astrophysical study cases, including the GeV–TeV emission from Markarian 501 and ultra-high-energy photons from a population of BL Lac objects.
While current observational data, such as the upper limits from the Pierre Auger Observatory, do not yet constrain the explored LIV parameters, our results demonstrate that astrophysical photon spectra provide a promising avenue for probing violations of Lorentz symmetry. Future multi-messenger observations, with improved sensitivity to both particle and photon fluxes, may offer critical tests of fundamental symmetries at the highest energies.
Acknowledgments
The authors are supported by the São Paulo Research Foundation (FAPESP) through grant number 2021/01089-1. VdS is supported by CNPq through grant number 308837/2023-1. MDF is supported by CAPES through grant number 88887.684414/2022-00. The authors acknowledge the National Laboratory for Scientific Computing (LNCC/MCTI, Brazil) for providing HPC resources for the SDumont supercomputer (http://sdumont.lncc.br).

### v5

@@
-Effects of Lorentz invariance violation on particle and photon production in astrophysical sources
+Effects of Lorentz invariance violation on charged particles and photon production in astrophysical sources
 Abstract
@@
 We derive modified expressions for radiated power and photon energy under LIV assumptions and incorporate them into first-order Fermi acceleration models.
-Our analysis reveals energy thresholds beyond which LIV significantly alters particle dynamics and photon spectra, introducing non-physical divergences that highlight limitations in perturbative approaches.
+Our analysis reveals energy thresholds beyond which LIV, within a kinematic framework, significantly alters particle dynamics and photon spectra, introducing non-physical divergences that highlight limitations in perturbative approaches.
 We model synchrotron self-Compton (SSC) emission in the presence of LIV and assess its consequences for photon fluxes from blazars, including Markarian 501 and the BL Lac population.
 LIV introduces distinct high-energy emission regions that deviate from standard expectations.
-Comparisons with observational data, particularly upper limits from the Pierre Auger Observatory, suggest that future multi-messenger observations could constrain LIV parameters through the non-detection of such excesses.
+Comparisons with observational data, particularly upper limits from the Pierre Auger Observatory, suggest that future multi-messenger observations, together with the full analysis of particle’s trajectories, could constrain LIV parameters through the non-detection of such excesses.
 1 Introduction
@@
 Nonetheless, a thorough investigation of the processes leading to the production of charged particles and photons under the hypothesis of LIV remains necessary [22].
-The mechanisms generating these particles and photons are fundamentally dependent on LI, which plays a critical role in determining the maximum energy, the emitted power, and the resulting energy spectrum [23].
+The mechanisms generating these charged particles and photons are fundamentally dependent on LI, which plays a critical role in determining the maximum energy, the emitted power, and the resulting energy spectrum [23].
 Consequently, such sources provide a natural framework for probing potential modifications of LI.
@@
 We also incorporate the LIV models for synchrotron and inverse Compton interactions into the context of first-order Fermi acceleration and photon emission in blazars.
-We investigate how the production of high-energetic particles and photons is affected by the LIV assumption.
+We investigate how the production of high-energetic charged particles and photons is affected by the LIV assumption.
 In Section 2, we present the LIV framework.
@@
 In this context, we independently evaluated the impact of LIV on both energy losses and energy spectrum of particles, as well as on the resulting photon emission.
-Section 5 presents a comparison between our theoretical predictions and observational data, including the identified blazar population and upper limits on photon flux from the Pierre Auger Observatory.
+Appendix A presents an illustrative comparison between our theoretical predictions and observational data, including the identified blazar population and upper limits on photon flux from the Pierre Auger Observatory.
 Finally, our conclusions and implications are summarized in Section 5.
@@
 $p=\frac{E}{\sqrt{1+\delta_{n}E^{n}}}\rm{.}$ (2.2)
-Lorentz Invariance Violation will also modify the Lorentz factor and, as shown in [35], this shall be done via the group velocity.
-Considering the definition $\gamma^{2}=\left(1-v^{2}\right)^{-1}$, as well as the group velocity, $v=\frac{\partial E}{\partial p}$, the LIV-modified Lorentz factor results in
-$\gamma^{2}_{\textrm{LIV}}=\frac{E^{2}}{{m^{2}-\left(n+1\right)\delta_{n}E^{n+2}}}\textrm{,}$ (2.3)
-where the high-energy approximation was used.
-As can be observed, the Lorentz factor increases even faster with energy under the LIV assumption and tends to infinity when $E^{n+2}\rightarrow\frac{m^{2}}{(n+1)\delta_{n}}$.
-Once this occurs, the Lorentz factor becomes negative, revealing a fundamental limitation of the current model and indicating the need for a non-perturbative treatment.
+No preferred reference frame is assumed, so the modified dispersion relation applies in every frame [35, 36, 37].
+Lorentz Invariance Violation can also lead to modifications in the Lorentz factor, which are naturally obtained by considering the group velocity.
+A related discussion in the context of deformed special relativity can be found in [38].
+To obtain this modification, one may start from the definition of the Lorentz factor, $\gamma^{2}=(1-v^{2})^{-1}$, which, when combined with the group velocity, $v=\partial E/\partial p$ [11, 4], leads to
+$\gamma^{2}=\frac{1}{1-(\frac{\partial E}{\partial p})^{2}}=\frac{E^{2}}{E^{2}-p^{2}-(n+2)\delta_{n}p^{n+2}-\left(\tfrac{n+2}{2}\right)^{2}(\delta_{n})^{2}p^{2n+2}}\ \textrm{.}$ (2.3)
+Recalling that $E^{2}-p^{2}=m^{2}+\delta_{n}p^{n+2}$, this expression can be written as
+$\gamma^{2}=\frac{E^{2}}{m^{2}-(n+1)\delta_{n}p^{n+2}-(\frac{n+2}{2})^{2}(\delta_{n})^{2}p^{2n+2}}\,\textrm{.}$ (2.4)
+It should be noted that this perturbative treatment may yield unphysical results for the modified factor (such as negative values or divergences).
+Therefore, we restrict our analysis to the parameter domain of $\delta_{n}$ in which the Lorentz factor remains physically meaningful.
+Moreover, the correction term $\delta_{n}p^{n+2}$ dominates over the quadratic contribution $(\delta_{n}p^{n+1})^{2}$ up to scales of $p^{n}\sim(\delta_{n})^{-1}$.
+However, in this extreme regime the Lorentz factor would already become negative, thus falling outside the physically acceptable domain.
+Consequently, we can safely neglect the quadratic term within the energy and $\delta_{n}$ ranges of interest.
+Under these considerations, and in the high-energy limit, we obtain the final expression for the Lorentz factor modified by LIV:
+$\gamma^{2}_{\textrm{LIV}}=\frac{E^{2}}{{m^{2}-\left(n+1\right)\delta_{n}E^{n+2}}}\textrm{.}$ (2.5)
+From equation 2.5, the allowed domain for the values of $\delta_{n}$ becomes evident.
+For each fixed pair of mass and energy, the LIV parameter is restricted to the region $\delta_{n}<\frac{m^{2}}{(n+1)E^{n+2}}$.
+For values outside this domain, the Lorentz factor diverges or becomes negative, revealing a fundamental limitation of the current model and indicating the need for a non-perturbative treatment.
 Different particles may exhibit different breaking parameters [4].
@@
 $E_{\gamma}=0.29\cdot\frac{3}{2}\gamma^{2}\frac{qB}{m}\,.$ (3.3)
-Using equation 3.4
+The previous equations can be used to analyze synchrotron emission when LIV is considered.
+Here we adopt a kinematic approach, in which LIV effects are included only through the modification of the Lorentz factor and the corresponding photon energy.
+This method does not alter the underlying quantum electrodynamics (QED) interaction vertices, and is intended as a first-order approximation to explore potential LIV effects on synchrotron emission arising exclusively from the modified dispersion relation.
+Therefore, using equations 3.2 and 3.3, along with 2.5:
 $P^{\text{LIV}}=\frac{2}{3}\frac{q^{4}}{m^{4}}|\vec{v}\times\vec{B}|^{2}\left[\frac{E^{2}}{1-\left(n+1\right)\frac{\delta_{n}}{m^{2}}E^{n+2}}\right]$ (3.4)
@@
 $E_{\gamma}^{\text{LIV}}=0.29\cdot\frac{3}{2}\cdot\frac{qB}{m^{3}}\left[\frac{E^{2}}{1-\left(n+1\right)\frac{\delta_{n}}{m^{2}}E^{n+2}}\right]\rm{.}$ (3.5)
-LIV has a significant effect on the synchrotron radiation.
-The modification of the Lorentz factor leads to a discontinuity at $E^{n+2}=\frac{m^{2}}{(n+1)\delta_{n}}$ in the power emitted and energy of the emitted photons.
-For particles with energy $E^{n+2}>\frac{m^{2}}{(n+1)\delta_{n}}$, the energy of the emitted photon becomes negative.
-These consequence of the calculations indicate a limitation in the perturbative approach used here.
-Introducing new physics for scales beyond the breaking point may be necessary to fully describe such events.
-Similar consequences were obtained by different approaches [39, 40].
+LIV modifies the Lorentz factor, which in turn affects the synchrotron power and characteristic photon energy.
+In the present kinematic framework, these equations capture the leading-order effects of LIV on the emission spectrum.
+For particles with energy $E^{n+2}>\frac{m^{2}}{(n+1)\delta_{n}}$, the perturbative approach breaks down, signaling that a full dynamical treatment, possibly including modified interaction vertices, would be required to fully describe such extreme events.
+However, as we are restricting the analysis to the physically allowed dominion of $\delta_{n}$, the present approach provides a reasonable first-order estimate of kinematic LIV effects on the synchrotron emission.
+Similar qualitative conclusions regarding the leading effect of LIV on synchrotron emission were obtained in full LIV-QED frameworks [42, 43].
 3.2 Inverse Compton under LIV assumption
@@
 where $p$ denotes the charged particle, $\gamma$ denotes photon and primed quantities refer to final states.
-The scattering cross section of relativistic particles is given by the Klein-Nishina formula [24]
-[…]
+Within our kinematic framework, the scattering cross section of relativistic particles is given by the Klein-Nishina formula [24]
+$\sigma_{\textrm{KN}}=\frac{r_{e}^{2}}{2}\int\left(\frac{E^{\prime}_{\gamma}}{E_{\gamma}}\right)^{2}\left(\frac{E^{\prime}_{\gamma}}{E_{\gamma}}+\frac{E_{\gamma}}{E^{\prime}_{\gamma}}-\sin^{2}{\theta}\right)d\Omega\textrm{,}$ (3.7)
 and the average scattering angle can be calculated through
-[…]
-Introducing LIV via the energy-momentum in Equation 2.1, the spatial components ($\mu=1,2,3$) become
+$\left<\cos{\theta}\right>=\frac{\int\cos{\theta}\frac{d\sigma}{d\Omega}d\Omega}{\int\frac{d\sigma}{d\Omega}d\Omega}\textrm{,}$ (3.8)
+such that a full dynamic treatment could introduce additional corrections to the cross-section [44, 10] and possibly alter the scattering angle distribution.
+As we restrict our analysis to a kinematic approach with no preferred reference frame, LIV can be introduced via the energy-momentum in Equation 2.1 in the charged particle frame of reference.
+With this modification, the spatial components ($\mu=1,2,3$) become
 $E^{{}^{\prime}2}_{p}=m^{2}+\left(\frac{E_{\gamma}^{{}^{\prime}2}}{1+\delta_{n}^{(\gamma)}E_{\gamma}^{{}^{\prime}n}}+\frac{E_{\gamma}^{2}}{1+\delta_{n}^{(\gamma)}E_{\gamma}^{n}}-\frac{2E_{\gamma}E_{\gamma}^{{}^{\prime}}\cos{\theta}}{\sqrt{\left(1+\delta_{n}^{(\gamma)}E_{\gamma}^{{}^{\prime}n}\right)\left(1+\delta_{n}^{(\gamma)}E_{\gamma}^{n}\right)}}\right)\times\\ \times\left[1+\delta_{n}\left(\frac{E_{\gamma}^{{}^{\prime}2}}{1+\delta_{n}^{(\gamma)}E_{\gamma}^{{}^{\prime}n}}+\frac{E_{\gamma}^{2}}{1+\delta_{n}^{(\gamma)}E_{\gamma}^{n}}-\frac{2E_{\gamma}E_{\gamma}^{{}^{\prime}}\cos{\theta}}{\sqrt{\left(1+\delta_{n}^{(\gamma)}E_{\gamma}^{{}^{\prime}n}\right)\left(1+\delta_{n}^{(\gamma)}E_{\gamma}^{n}\right)}}\right)^{n/2}\right]^{1/2}\textrm{,}$ (3.9)
@@
 However, here, we present a generalization for all values of $n$.
-For this interaction, LIV effects could be considered in the charged particle sector and the photon sector.
-In the following, we focus on the photon sector and defer the full treatment of the charged-particle sector to future work.
+For this interaction, LIV effects could, in principle, appear in both the photon and charged-particle sectors.
+In this work, we focus on the photon sector, where the modified dispersion relation directly affects the kinematics of the scattering process.
+Although several studies have derived strong bounds on LIV coefficients for charged particles, mainly from the non-observation of vacuum Cherenkov radiation [46, 13, 47, 48], such analyses typically treat each sector independently and neglect combined propagation and cascade effects, a methodological simplification that we also adopt here to isolate the photon contribution.
 We set $\delta_{n}=0$ to isolate effects on the photon sector and solved Equation 3.10 numerically.
@@
 Figure 2 shows the evolution of the cross section with the initial energy of the photon, including LIV via Equations 3.9 and 3.10.
-LIV significantly impacts $\sigma_{\textrm{KN}}$, enhancing the interaction probability and effectively bypassing Klein-Nishina suppression until the Thomson limit is recovered.
+In our kinematic approach, LIV modifications to the photon dispersion relation lead to changes in the effective cross section, enhancing the interaction probability and effectively bypassing Klein-Nishina suppression until the Thomson limit is recovered.
 Figure 3 shows the evolution of the scattering angle with the initial energy of the photon including LIV. A new transition to the Thomson regime is observed at high energies.
 Using the modified cross section and final energy of the photon, the emitted power under LIV assumption can be calculated via
-[…]
-where $\eta(E_{\gamma})$ is the energy distribution of photons before the interaction. LIV increases both the final energy and cross section, thus amplifying the emitted power.
+$P=\int\sigma_{\textrm{KN}}\left[E^{\prime}_{\gamma}-E_{\gamma}\right]\eta(E_{\gamma})dE_{\gamma}\textrm{,}$ (3.11)
+where $\eta(E_{\gamma})$ is the energy distribution of photons before the interaction. Within our framework, LIV increases both the final energy and cross section, thus amplifying the emitted power.
+As mentioned in Section 2, we do not introduce background timelike vectors or coefficients that would single out a preferred reference frame.
+Our analysis thus focuses exclusively on the modifications arising from the photon dispersion relation, without assuming any model-specific preferred-frame structure.
+Nonetheless, certain LIV formulations do predict the existence of a preferred frame, which could alter the mapping of scattering angles and change the average scattering angle obtained.
+As shown in Figure 3, angular corrections would be expected to appear at comparatively lower energies, while LIV modifications to the boost from the preferred frame to the charged particle frame are expected to become relevant only when $E^{n}\sim\delta_{n}^{-1}$.
+The energy scale at which the Thomson-like regime is recovered in our kinematic analysis coincides with the scale where LIV-induced boost effects would likely emerge.
+Hence, the actual existence and properties of this new regime remain uncertain until a complete analysis is performed within a specific LIV model that consistently includes both kinematic and dynamical effects.
+A full LIV-QED framework could introduce further corrections to the scattering amplitude and radiative processes.
+Nevertheless, the present treatment provides a controlled first step, highlighting the dominant kinematic effects within the inverse Compton process and offering a useful benchmark for comparison with more complete approaches.
 4 First-order Fermi acceleration including radiation process under LIV assumption
@@
 In a previous publication [22], we investigated how LIV modifies the energy gain per scattering, the acceleration timescales, and the spectral indices of particles accelerated by the first-order Fermi mechanism.
-In this Section, we will merge the first-order Fermi with synchrotron and SSC energy losses, both including LIV.
+In this Section, we will merge the first-order Fermi with synchrotron and SSC energy losses, both including LIV within our kinematic framework.
+It is important to note that, although we adopt a superluminal LIV framework, the restrictions imposed on the parameter $\delta_{n}$, as mentioned in Section 2, ensures that particle velocities approach, but never exceed, the speed of light.
+Consequently, we do not consider other energy losses from processes that would otherwise become kinematically allowed, such as vacuum Cherenkov emission.
 4.1 Synchrotron radiation in first-order Fermi acceleration under LIV assumption
@@
 In Reference [22], we calculated the energy gained by particles under first-order Fermi acceleration considering LIV.
-This modified energy gain, along with the power lost through synchrotron radiation, gives the following net gain $-b$:
-$b=-\frac{\left<\Delta E\right>}{\tau_{avg}}+P_{sync}\frac{\tau_{avg}}{\tau_{avg}}\rm{,}$ (4.1)
+This modified energy gain, along with the power lost through synchrotron radiation, gives the following net gain b-b over time:
+$b\cdot\tau_{avg}=-\left<\Delta E\right>+P_{sync}\cdot\tau_{avg}\rm{,}$ (4.1)
 where $\tau_{avg}$ is the average crossing time for the particle traversing the shock wave. We consider the best-case scenario of Bohm diffusion [52]
 $\tau_{avg}\approx\frac{5}{qBV}\frac{E}{\sqrt{1+\delta_{n}E^{n}}}\rm{.}$ (4.2)
-In a scenario where synchrotron radiation is the dominant energy loss during shock wave acceleration, we can use Equation 3.4 to determine the maximum energy a charged particle can gain through first-order Fermi acceleration [46] as the equilibrium between loss and energy gain, $\Delta E\right>=P_{sync}\cdot\tau_{avg}$, which gives
+In a scenario where synchrotron radiation is the dominant energy loss during shock wave acceleration, we can use Equation 3.4 to determine the maximum energy a charged particle can gain through first-order Fermi acceleration [53] as the equilibrium between loss and energy gain, $\left<\Delta E\right>=P_{sync}\cdot\tau_{avg}$, which gives
 $\frac{E_{\textrm{max}}^{2}}{\left[1-\left(n+1\right)\frac{\delta_{n}}{m^{2}}E^{n+2}\right]}\approx\frac{m^{4}}{B}\rm{.}$ (4.3)
@@
 A transition between the two regimes is very clear.
-In the regime of low $\delta_{n}$ values, the maximum energy is determined by the balance between energy gain and synchrotron losses.
-As $\delta_{n}$ increases, LIV becomes more significant, the Lorentz factor barrier dominates, such that a particle reaching $E^{n+2}=\frac{m^{2}}{(n+1)\delta_{n}}$ would emit all its energy, as shown in Equation 3.4.
-This transition is shown as an abrupt energy loss in these plots.
-4.1.2 Synchrotron: energy spectrum of particles and photons
+In the regime of low $\delta_{n}$ values, the maximum energy is determined by the balance between energy gain and synchrotron losses for standard LI.
+As $\delta_{n}$ increases, kinematic LIV becomes more significant, and as $E^{n+2}\rightarrow\frac{m^{2}}{(n+1)\delta_{n}}$ the emitted power rises sharply, constraining the maximum attainable energy.
+This transition appears as an abrupt energy loss in the plots.
+4.1.2 Synchrotron: energy spectrum of charged particles and photons
 We will use the one-zone model to calculate the energy spectrum of particles accelerated in the source [54].
 This model considers particles traveling through an acceleration and a cooling (or energy-loss) region.
-The transport of particles can be calculated using the diffusion-loss equation to describe the number of particles $N$ within an energy range $E$ to $E+\delta E$ at a given time
+The transport of particles can be calculated using the diffusion-loss equation to describe the number of particles $N$ within an energy range $E$ to $E+dE$ at a given time
 $\frac{dN}{dt}+\frac{d}{dE}\left[b\cdot N\right]=-\frac{N}{\tau_{scp}}+Q_{\mathrm{inj}}\rm{,}$ (4.4)
@@
 A magnetic field of $B=10^{-10}\ \textrm{T}$ was used.
-The suppression in the energy spectrum is a direct consequence of the barrier created by the Lorentz factor.
-As shown in section 4.1.1, above a certain maximum energy, which depends on the LIV break, energy losses are dominant, and particles stop gaining energy.
+The abrupt decay in the spectrum is a direct consequence of the maximum energy derived in Section 4.1.1.
+Although the emitted power formally diverges at higher energies, in our analysis we restrict ourselves to values of energy and $\delta_{n}$ for which the Lorentz factor remains physically valid.
+In this sense, the behavior observed in the plots reflects the trend expected when kinematic LIV is taken into account.
+To obtain quantitative bounds from this behavior, however, a complete dynamical treatment would be required given that the divergences and unphysical results could be modified in a fully consistent framework.
 The energy spectrum of photons produced by the accelerated particles via synchrotron emission can be modeled by [23]
-[…]
+$N_{\gamma}(E_{\gamma})=\frac{D}{4\pi}\int_{E_{min}}^{E_{max}}P(E)N_{p}(E)dE\rm{,}$ (4.5)
 where $P$ is the synchrotron power, $N_{p}$ is the particle’s energy spectrum, and $D$ is the volume of the source.
@@
 Figures 8 and 9 show the resulting energy spectrum for electrons and protons, respectively. In these calculations, we considered a source with a size of $1\hskip 3.00003pt\rm{pc}$.
-LIV leads to the emergence of a distinct high-energy emission, where the final photon energy exhibits rapid growth near the divergence threshold - $E^{n+2}\rightarrow\frac{m^{2}}{(n+1)\delta_{n}}$ expected from Equation 3.5.
+LIV leads to the emergence of a distinct high-energy emission, where the final photon energy exhibits rapid growth near the divergence threshold - $E^{n+2}\rightarrow\frac{m^{2}}{(n+1)\delta_{n}}$ - expected from Equation 3.5.
 Notably, this region contributes a non-negligible flux, consistent with the increasing emitted power derived in Equation 3.4.
@@
 The energy range where this violation occurs is indicated by a dashed line in Figures 8 and 9.
-Lorentz invariance violation not only introduces an unanticipated energy domain in synchrotron emission but could also alter SSC spectra through the inclusion of higher-energy photons in the radiation background, resulting in emissions at previously unexpected energies and potentially altering the predicted spectra.
-These modifications could serve as a means to constrain LIV, as the non-detection of such new energy regions could impose limits on the LIV parameters.
+As noted earlier, these results were obtained solely by incorporating the modified dispersion relation, with the goal of isolating the dominant kinematical LIV effects on synchrotron emission.
+A full analysis of this process within a dynamical LIV–QED framework may confirm or modify these conclusions, and we leave such an investigation for future work.
+Should the current trend persist, LIV not only introduces an unanticipated energy domain in synchrotron emission but could also alter SSC spectra through the inclusion of higher-energy photons in the radiation background, resulting in emissions at previously unexpected energies and potentially altering the predicted spectra.
+These modifications could serve as a means to constrain LIV, especially when combined with propagation effects, since additional processes such as photon splitting [47, 17] are expected to become relevant in this extended energy range of emission.
 4.2 Synchrotron self-Compton in first-order Fermi acceleration under LIV assumption
@@
 Notice that $\delta_{n}^{(\gamma)}$ will not affect the photon distribution; only the inverse Compton scattering will be affected by this parameter.
-Again, the parameter $\delta_{n}$ will be set to 00 in order to focus exclusively on the behavior of LIV in photons.
-4.2.1 Synchrotron self-Compton: energy spectrum of particles and photons
+Again, the parameter $\delta_{n}$ will be set to 0 in order to focus exclusively on the behavior of LIV in photons.
+4.2.1 Synchrotron self-Compton: energy spectrum of charged particles and photons
 In this section, we repeat the calculations in Subsection 4.1.2, now considering SSC as the dominant energy-loss mechanism.
@@
 In contrast, for systems dominated by synchrotron losses, only emission modifications persist without spectral suppression, as shown in Figure 14.
-5 Radiation emission from blazars
-In the previous sections, idealized cases of acceleration, emission, and related processes were considered to highlight the new LIV phenomena introduced in each case.
-In this section, we explore a more realistic scenario of photon production including LIV in Markarian 501 and in blazars.
-5.1 GeV-TeV photon flux from Markarian 501
-We used the models developed in the previous sections to calculate the GeV-TeV photon flux generated in Mkr 501.
-We used the data reference [50], in particular, the magnetic field strength and the electron energy spectrum responsible for both synchrotron emission and the inverse Compton process.
-We assume a cutoff at $10^{14}\ \rm{eV}$, consistent with the maximum energy predicted by Equation 4.3 under the magnetic field conditions of Mrk 501.
-Figure 15 shows the energy spectrum of photons for two values of LIV break when synchrotron emission is considered according to section 3.1.
-The emission extends to higher energies, effectively bypassing the cutoff imposed by the LI predictions.
-Additionally, Figure 16, which presents a magnified view of the synchrotron peak, shows that for sufficiently large LIV parameters, even the low-energy portion of the spectrum may be altered, leading to a noticeable divergence from the observational data.
-Figure 17 shows the energy spectrum of photons for two values of LIV break when SSC emission is considered according to section 3.2.
-A second emission region is predicted under LIV assumption which is not present in LI models.
-This new component in the flux allows one to open a new window to constrain the parameters associated with LIV.
-The non-detection of the predicted high-energy photons would impose upper limits on the extent of the violation.
-5.2 Ultra-high energy photons from BL Lac
-The new emission regions arising from synchrotron and SSC processes also populate the photon spectrum at energies above $10^{17}\ \rm{eV}$.
-To illustrate this possibility, we compare the predicted LIV-induced emissions with the upper limits on the photon flux established by the Pierre Auger Observatory [51, 52].
-The flux is modeled as originating from a population of BL Lac objects, given that approximately $10^{3}$ such sources have been identified [53, 54].
-For simplicity, the combined flux is computed by summing the contributions of $10^{3}$ identical sources without accounting for source distribution, propagation effects, or intrinsic source-to-source variations, and adopting typical parameters for BL Lac objects [55].
-We note that the determination of average source parameters typically does not aim to constrain the cutoff energy of the electron energy spectrum, as its impact on the overall source characteristics is minimal.
-Therefore, in this work we adopt a cutoff energy of $10^{14}\ \rm{eV}$.
-Figures 18 and 19 show the energy spectrum of synchrotron photons calculated with this simplified model considering LIV for $n=1$ and $n=2$, respectively.
-We show the upper limits on the photon flux measured by the Pierre Auger Observatory for comparison in Figures 20 and 21.
-A similar procedure was applied to the SSC emission considering LIV.
-The energy spectrum of photons are presented in Figures 22 and 23 for $n=1$ and $n=2$, respectively.
-Additionally, we consider a scenario without an intrinsic cutoff in the electron energy distribution, illustrating the most optimistic case for high-energy photon emission.
-These comparisons between the flux and the Pierre Auger Observatory limits can be seen in Figures 24 and 25.
-At present, the comparison with data from the Pierre Auger Observatory does not yield constraints on the investigated LIV parameters.
-Establishing meaningful limits would require a more detailed astrophysical treatment, including the effects of source distribution, cosmological propagation, and potential contributions from other classes of sources, such as flat-spectrum radio quasars.
-Increasing the number of detected sources would enhance the cumulative photon flux, and a detailed analysis incorporating these effects could potentially lead to significant constraints on LIV.
+The same discussion presented in Subsection 4.1.2 regarding the emergence of new emission regimes and their impact on the particle spectrum also applies here.
+We highlight the exploratory character of these results and outline a possible strategy for incorporating LIV effects into the process.
 5 Conclusion
-In this work, we examined the influence of Lorentz invariance violation (LIV) on the production and emission of high-energy particles and photons in astrophysical environments.
+In this work, we examined the influence of Lorentz invariance violation (LIV) on the production and emission of high-energy charged particles and photons in astrophysical environments.
 By modifying synchrotron and inverse Compton processes within the LIV framework, we derived the resulting changes in radiated power, photon energy, and energy spectra of accelerated particles.
 Our results demonstrate that LIV induces rapid growth in the Lorentz factor at high energies, leading to divergences in emission characteristics and suggesting a breakdown of the perturbative formalism, requiring new theoretical treatments at extreme energies.
+The model employed here is purely kinematical; a fully dynamical analysis could introduce additional corrections to both cross sections and emission spectra, and may alter some of the behaviors reported in this work.
 Incorporating these effects into first-order Fermi acceleration models and synchrotron self-Compton scenarios, we showed that LIV modifies both particle spectra and the resulting multi-wavelength photon emission.
-In particular, LIV predicts the emergence of distinct high-energy components in the photon spectrum that are absent under Lorentz-invariant assumptions.
-These features were evaluated in astrophysical study cases, including the GeV–TeV emission from Markarian 501 and ultra-high-energy photons from a population of BL Lac objects.
-While current observational data, such as the upper limits from the Pierre Auger Observatory, do not yet constrain the explored LIV parameters, our results demonstrate that astrophysical photon spectra provide a promising avenue for probing violations of Lorentz symmetry.
-Future multi-messenger observations, with improved sensitivity to both particle and photon fluxes, may offer critical tests of fundamental symmetries at the highest energies.
+In particular, LIV predicts the emergence of distinct high-energy components in the photon spectrum, while suppressing the energy spectrum of charged particles.
+These features of the photon emission were evaluated in astrophysical study cases, including the GeV–TeV emission from Markarian 501 and ultra-high-energy photons from a population of BL Lac objects (Appendix A).
+Looking forward, once a consistent dynamical formulation of LIV is established to ensure the validity of these behaviors, a multi-messenger analysis becomes a natural extension.
+LIV could simultaneously impact the photon sector and the spectrum of charged particles accelerated through first-order Fermi mechanisms, establishing correlated signatures across different messengers.
+While current observational data, such as the upper limits from the Pierre Auger Observatory, do not yet constrain the explored LIV parameters, our results demonstrate that astrophysical photon spectra provide a promising avenue for probing violations of Lorentz symmetry, especially if the results here presented are maintained in a fully consistent dynamical formulation.
+Future multi-messenger observations, with improved sensitivity to both charged particle and photon fluxes, may offer critical tests of fundamental symmetries at the highest energies.
 Acknowledgments
@@
 The authors acknowledge the National Laboratory for Scientific Computing (LNCC/MCTI, Brazil) for providing HPC resources for the SDumont supercomputer (http://sdumont.lncc.br).
+Appendix A Illustrative application to blazars
+In the main text, we analyzed idealized cases of acceleration, emission, and related processes in order to isolate the effects introduced by Lorentz invariance violation (LIV).
+As an illustrative extension, we present here a simplified application to blazars, focusing on the case of Markarian 501 and, more generally, on the BL Lac population.
+The purpose of this appendix is not to derive quantitative constraints on the LIV coefficients, but rather to highlight possible observational signatures in scenarios where dynamical effects and the existence of preferred frames do not alter the trends previously reported.
+Using the framework developed in the main text, we applied the synchrotron and synchrotron self-Compton (SSC) models to the case of Mrk 501, adopting representative source parameters from the literature [57].
+The resulting synchrotron spectra, shown in Figures 15 and 16, indicate that LIV effects can extend the emission beyond the standard LI cutoff and even modify the low-energy part of the spectrum for sufficiently large LIV parameters.
+In the SSC channel, the predicted spectra displayed in Figure 17 reveal the emergence of an additional high-energy emission region that is absent in LI models.
+Extending the same procedure to a simplified population of BL Lac objects, we modeled each source using typical parameters reported in Ref.[58], and then computed the cumulative photon spectrum as the simple sum of contributions from approximately $10^{3}$ known BL Lac objects [59, 60].
+This cumulative photon spectrum reaches energies above $10^{17}\ \rm{eV}$, as illustrated in Figures 1819 for synchrotron emission and in Figures 2223 for SSC emission.
+For reference, these fluxes are also compared with the upper limits on the ultra-high-energy photon flux reported by the Pierre Auger Observatory in Figures 2021 (synchrotron) and Figures 2425 (SSC).
+All fluxes shown were computed in the vicinity of the sources, so propagation effects were not taken into account.
+We stress, however, that these results are only meant as an illustration.
+Deriving robust limits on LIV coefficients from this type of analysis would require a series of additional steps, including: (i) implementing a full dynamical treatment of synchrotron and inverse Compton processes, rather than a purely kinematical one; (ii) modeling photon propagation with LIV, including absorption in the extragalactic background light and possible modifications to pair production thresholds; (iii) accounting for the development of atmospheric showers at ultra-high energies, which directly impacts observational signatures; (iv) incorporating realistic source populations with distributions in distance, luminosity, and variability; and (v) considering other relevant source classes such as flat-spectrum radio quasars.
+Each of these ingredients can alter the resulting spectrum in a nontrivial way and must be included before quantitative bounds can be established.
+Within these limitations, the present appendix should be regarded as a first step, illustrating the qualitative trends expected when LIV is incorporated into blazar emission models.
+A complete treatment combining dynamics, propagation, and population effects will be required in future work to transform these signatures into quantitative constraints on LIV parameters.
+Figure 1: Final photon energy in the reference frame of the electron as a function of its initial energy.
+A scattering angle of $\pi/2$ was adopted.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for the first two orders of violation.
+The electron LIV parameter was set to 0.
+The result for the first violation order was previously obtained in [45].
+Figure 2: Klein-Nishina cross section for a photon-electron interaction.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for the first two orders of violation.
+The electron LIV parameter was set to 0.
+The result for the first violation order was previously obtained in [45].
+Figure 3: Average of the cosine of the scattering angle of photons as a function of the initial photon energy.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for the first two orders of violation.
+The electron LIV parameter was set to 0.
+Figure 4: Maximum electron energy as a function of breaking parameter $\delta_{n}^{(e)}$.
+Electrons are accelerated via the first-order Fermi mechanism and are susceptible to synchrotron losses.
+A magnetic field of $B=10^{-10}\ T$ was adopted.
+Both cases of $n=1$ and $n=2$ are shown.
+Figure 5: Maximum proton energy as a function of the breaking parameter $\delta_{n}^{(p)}$.
+Protons are accelerated via the first-order Fermi mechanism and are susceptible to synchrotron losses.
+A magnetic field of $B=10^{-10}\ T$ was adopted.
+Both cases of $n=1$ and $n=2$ are shown.
+Figure 6: Electron’s energy spectrum with an arbitrary normalization.
+The black solid line represents the Lorentz invariant scenario.
+The energy of the spectral index break was taken to be $E_{b}=10^{13}\ \rm{eV}$.
+Different values of $\delta_{p}^{(e)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+For each fixed value of $\delta_{n}$, the energy of the electron is restricted to $E^{n+2}<\frac{m^{2}}{(n+1)\delta_{n}}$.
+Figure 7: Proton’s energy spectrum with an arbitrary normalization.
+The black solid line represents the Lorentz invariant scenario.
+The energy of the spectral index break was taken to be $E_{b}=10^{20}\ \rm{eV}$.
+Different values of $\delta_{n}^{(p)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+For each fixed value of $\delta_{n}$, the energy of the proton is restricted to $E^{n+2}<\frac{m^{2}}{(n+1)\delta_{n}}$.
+Figure 8: Photon’s energy spectrum emitted via synchrotron emission by a population of electrons with an arbitrary normalization.
+The energy of the spectral index break was taken to be $E_{b}=10^{13}\ \rm{eV}$.
+A magnetic field of $B=10^{-10}\ T$ was adopted.
+The black solid line represents the Lorentz invariant scenario.
+The dashed lines indicate the energy region at which energy conservation is violated.
+Different values of $\delta_{n}^{(e)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+For each fixed value of $\delta_{n}$, the energy of the electron emitting radiation is restricted to $E^{n+2}<\frac{m^{2}}{(n+1)\delta_{n}}$.
+Figure 9: Photon’s energy spectrum emitted via synchrotron emission by a population of protons with an arbitrary normalization.
+The energy of the spectral index break was taken to be $E_{b}=10^{20}\ \rm{eV}$.
+A magnetic field of $B=10^{-10}\ T$ was adopted.
+The black solid line represents the Lorentz invariant scenario.
+The dashed lines indicate the energy region at which energy conservation is violated.
+Different values of $\delta_{n}^{(p)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+For each fixed value of $\delta_{n}$, the energy of the electron emitting radiation is restricted to $E^{n+2}<\frac{m^{2}}{(n+1)\delta_{n}}$.
+Figure 10: Electron’s energy spectrum resulting from the one-zone model with an arbitrary normalization.
+The electron breaking parameter was set to 0.
+The energy of the spectral index break was taken to be $E_{b}=10^{13}\ \rm{eV}$.
+The black solid line represents the Lorentz invariant scenario.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+Figure 11: Proton’s energy spectrum resulting from the one-zone model with an arbitrary normalization.
+The proton breaking parameter was set to 0.
+The energy of the spectral index break was taken to be $E_{b}=10^{20}\ \rm{eV}$.
+The black solid line represents the Lorentz invariant scenario.
+The dashed lines indicate the energy region at which energy conservation is violated.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+Figure 12: Photon’s energy spectrum emitted via synchrotron self-Compton by a population of electrons with an arbitrary normalization.
+Inverse Compton radiation was considered the dominant loss process.
+The energy of the spectral index break was taken to be $E_{b}=10^{11}\ \rm{eV}$.
+The black solid line represents the Lorentz invariant scenario.
+The dashed lines indicate the energy region at which energy conservation is violated.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+Figure 13: Photon’s energy spectrum emitted via synchrotron self-Compton by a population of protons with an arbitrary normalization.
+Inverse Compton radiation was considered the dominant loss process.
+The energy of the spectral index break was taken to be $E_{b}=10^{20}\ \rm{eV}$.
+The black solid line represents the Lorentz invariant scenario.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+Figure 14: Photon’s energy spectrum emitted via synchrotron self-Compton by a population of electrons with an arbitrary normalization.
+Synchrotron emission was considered the dominant loss process.
+The energy of the spectral index break was taken to be $E_{b}=10^{11}\ \rm{eV}$.
+The black solid line represents the Lorentz invariant scenario.
+The dashed lines indicate the energy region at which energy conservation is violated.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for the studied orders of violation, $n=1$ and $n=2$.
+Figure 15: Photon’s energy spectrum resulting from the synchrotron emission of Markarian 501.
+The data was taken from [19] and only the synchrotron peak is shown.
+Different values of $\delta_{n}^{(e)}$ are shown for both violation orders, $n=1$ and $n=2$.
+Figure 16: Zoom on the photon’s energy spectrum resulting from the synchrotron emission of Markarian 501.
+The data was taken from [19] and only the synchrotron peak is shown.
+Different values of $\delta_{n}^{(e)}$ are shown for both violation orders, $n=1$ and $n=2$.
+Figure 17: Photon’s energy spectrum resulting from the SSC emission of Markarian 501.
+The data was taken from [19] and only the SSC peak is shown.
+Different values of $\delta_{n}^{(\gamma)}$ are shown for both violation orders, $n=1$ and $n=2$.
+Figure 18: Photon’s energy spectrum resulting from the synchrotron emission of a BL Lac object.
+A cut-off of $10^{14}\ \rm{eV}$ was adopted.
+The first order of violation, $n=1$, is shown for different breaking parameters in electrons.
+Figure 19: Photon’s energy spectrum resulting from the synchrotron emission of a BL Lac object.
+A cut-off of $10^{14}\ \rm{eV}$ was adopted.
+The second order of violation, $n=2$, is shown for different breaking parameters in electrons.
+Figure 20: Integral photon flux for energies beyond $10^{17}\ \rm{eV}$ for a violation of order $n=1$ adopting different values of $\delta_{1}^{(e)}$.
+Red dots are the current upper limits set by the Pierre Auger Observatory.
+Figure 21: Integral photon flux for energies beyond $10^{17}\ \rm{eV}$ for a violation of order $n=2$ adopting different values of $\delta_{2}^{(e)}$.
+Red dots are the current upper limits set by the Pierre Auger Observatory.
+Figure 22: Photon’s energy spectrum resulting from the SSC emission of a BL Lac object.
+A cut-off of $10^{14}\ \rm{eV}$ was adopted.
+The first order of violation, $n=1$, is shown for different breaking parameters in photons.
+Figure 23: Photon’s energy spectrum resulting from the SSC emission of a BL Lac object.
+A cut-off of $10^{14}\ \rm{eV}$ was adopted.
+The second order of violation, $n=2$, is shown for different breaking parameters in photons.
+Figure 24: Integral photon flux for energies beyond $10^{17}\ \rm{eV}$ for a violation of order $n=1$ adopting different values of $\delta_{1}^{(\gamma)}$.
+The solid curves are the integral of photon fluxes for a population of $10^{3}$ BL Lac sources with a cut-off of $10^{14}\ \rm{eV}$.
+Dashed lines represent the same emission without cut-offs.
+Red dots are the current upper limits set by the Pierre Auger Observatory.
+Figure 25: Integral photon flux for energies beyond $10^{17}\ \rm{eV}$ for a violation of order $n=2$ adopting different values of $\delta_{2}^{(\gamma)}$.
+The solid curves are the integral of photon fluxes for a population of $10^{3}$ BL Lac sources with a cut-off of $10^{14}\ \rm{eV}$.
+Dashed lines represent the same emission without cut-offs.
+Red dots are the current upper limits set by the Pierre Auger Observatory.